# The Magnetic Filling in MAD Simulations and its Impact on the Jet in M87

### v1

@@ 3.3 Jet Edge and Width Profiles
Figure 13: $\langle I\rangle=5\times 10^{-6}\,[\mathrm{Jy}]$ contours of the time averaged 86 GHz ray-tracings (not the line integral convolved images) shown in Fig. 11 for simulations K.80 (green) K.90 (blue) and K.99 (red). The colored areas are $1\sigma$-bands of the standard deviations in the toroidal direction.
Figure 14: Jet width of M87 at 86 GHz along the longitudinal jet axis for simulations K.80 (green) K.90 (blue) and K.99 (red) with corresponding $1\sigma$-bands. The black dots are data points of the jet width of M87 from the observation of Lu et al. (2023). The dotted line is a power-law fit from observations of M87 (source). The solid horizontal line is the estimated 86 GHz ring size and the dashed horizontal line is the 230 GHz ring size with respective error band (gray areas). The vertical lines are the distances from the BH, where the intrinsic half-opening angles equals the jet viewing angle of $\theta_{\mathrm{v}}=17^{\circ}$ (color-code is consistent).
From these jet profiles of the $86$ GHz ray-tracing we can now extract the jet widths as a function $W(z)$ of the jet longitudinal axis $z$ (as defined before) and compare these to the observations of Lu et al. (2023) shown in Fig. 14. We note that the data-points of the jet-width measurements do not show the error-bars in Lu et al. (2023), hence we cannot elaborate on the statistical certainty of this comparison. From Fig. 14, it is apparent that the theoretical predictions especially of K.80 and K.90 do match the observations well at the scales below $100\,\mu as$, contrarily to the claims of Lu et al. (2023) that MAD simulations may have difficulty explaining the observed jet width at these scales. At the scales shown above $100\,\mathrm{\mu as}$ the simulations K.80 and K.90 converge to the extrapolated $W(z)\propto z^{0.58}$ power law fit from large scale observations (see Asada & Nakamura (2012), Hada et al. (2013), Hada et al. (2016)). K.99 follows the power-law further up-stream and exhibits a sudden increase of the jet width at $\sim 40\,\mathrm{\mu as}$, whilst even closer to the BH it becomes almost constant, but even wider than the other models. This sudden increase in the width of the jet can be explained by the counter-jet, which exhibits a wider opening angle and thus shifts the jet edge of the counter-jet to positive $z$-values up to $\sim 100\,\mathrm{M}$ (cf. Fig. 11, right image). For the common choice of the R-$\beta$ electron temperature- and $\kappa$ non-thermal electron prescriptions chosen here K.99 exhibits a steeper slope than this observation suggests, rather following the large-scale power-law. The new observational data points at the scales closer to the jet launching region suggest that K.90 satisfies these observations very well, as the data points are very well within the variability of K.90 (cf. green band in Fig. 14). We also see that this holds mostly for the common model K.80, as well.
@@ 4 Conclusions
We find that the initial magnetic energy content of the torus systematically alters the jet energetics and morphology of the developed simulations. It affects mostly the EM contribution in the Jet and from that gains $\sim 60\%$ efficiency in the conversion from rest-mass energy to the energy of the outflow by means of increasing the magnetic energy of the torus, at the initialization. This in turn, can make nonthermal particle acceleration from vaccuum gaps and subsequent pair cascades more likely, due to the increased EM-field in the jet spine (cf. Levinson & Rieger 2011; Wendel et al. 2021, for details on the vacuum gap model and applications to low-luminosity AGN like M87 and Mrk501). We also find, that an increased magnetic filling of the FM torus elevates the outward angular momentum transport, primarily in the jet, driven by maxwell-stresses. Comparing the angular momentum tranport of BHs with very high spin presented here to Chatterjee & Narayan (2022), in which the case of a non-rotating BH in a standard MAD-setup ($F=0.8$) is investigated, the major differences are: $i)$ the total radial angular momentum flux is dominated by radially outward directed flux in the jet and wind, majorly originating from the jet base. ii)ii) Morpholigically, the angular momentum vector-fields are distinct, in that the high spin BH creates field-lines that point radially outwards in a polar angle range from $\sim(0-70)^{\circ}$ or $\sim(110-180)^{\circ}$. Towards the equatorial plane the field-lines transition from a monopol-geometry to a quadrupol-geometry in the disk and torus. The dominant, Maxwell-stress induced, outward polar flux is redirected to an equatorially inward, advectively driven flux. In contrast, the case of a non-rotating BH exhibits dominant advectively driven inward angular momentum flux in a larger portion of the disk and torus, whereas the outward directed maxwell-stress induced flux is subdominant in the jet, such that the net angular momentum is transferred to the central BH. The outward angular momentum transport also covers a much smaller polar angle compared to simulations with high spin, contributing far less to the strength of outward winds. Our simulations demonstrate the effect of extracting the angular momentum of the BH to the angular momentum of the outflowing plasma in the jet and wind as proposed by Blandford & Znajek (1977) for the first time in MADs with high spin BHs. We further find that there is a maximum of all net angular momentum fluxes inside the ergopsheric region of the BH, opposed to the net-constant fluxes for MAD-simulations with no spin.
@@
It appears that the disk-vertical flux tubes created by the magnetically arrested disk affect the jet structure more pronounced for high filling factors. From our sampled MAD states, we find a tendency for spatially more extended flux tubes that affect the jet morphology by introducing significant azimuthal asymmetries in the case of K.99 that propagate along the jet at least up to $200\,\mathrm{M}$. The jet sheath structure is more stable on these scales, i.e. the jet structure exhibits lower variability in the jet-perpendicular direction. However, since we restricted the jet morphology analysis to a time-interval, in that a single flux eruption occurs in each simulation, further detailed, statistical analysis of the impact of the filling factor on the quasi-periodic flux eruptions could provide valuable insights into the observable footprints that they have.
@@
The radiative transport calculations as well show that the jet-sheath structure is more stable and collimated up to $\sim 200\,\mu as$ supported by the higher EM energy flux contribution in K.99. We find significantly elevated LP fractions up to $\sim 15\%$ in agreement with observations of the EHT Collaboration et al. (2021b). The jet width relation over the jet axial distance shows that GRRT calculations of GRMHD simulations of magnetically arrested accretion flow is able to match the observed jet width profile of Lu et al. (2023) very well for the reference models. For K.99 we find deviations from the jet width, as it follows the observed large-scale power-law relation $W\propto z^{0.58}$ to shorter distances, whereas the observations follow a sub-power-law relation. The deviation of K.99 from the observations, however, is a result of the different techniques used for the determination of the jet width. In the observations of Lu et al. (2023) the HWHM of the outermost compenents perpendicular to the jet longitudinal axis determines the jet width, whereas we do not recreate the edge brightend jet morphology and thus use brightness contours to determine the jet width. This makes the detailed jet width relation difficult to compare to these observations, nevertheless we show that all our MAD simulations provide a wide jet base, substantially wider than the pure Blandford-Znajek (BZ) jet (cf. Blandford & Znajek 1977) for a high spin BH ($a\rightarrow 1$). This is evident from our analysis of the jet energy fluxes, that the almost evacuated, highly magnetized jet spine ($\sigma\gtrsim 1$) is dominated by EM energy flux and in the enveloping jet sheath the wind contributes to the outflowing kinetic energy flux increasing with distance. This jet sheath region ($\sigma\lesssim 1$ and the gravitationally unbound part of the plasma) is the origin of the jet emission and the $\sigma=1$ contour, hence, is not a good tracer for the outermost jet edge that could be observed. We remind here, that $\sigma=\sigma_{\mathrm{cut}}=1$ is the value for the emission cut-off for the jet spine, where we do not expect significant synchrotron emission in the radio-regime because of the extremly low electron densities, as well as efficient synchrotron cooling, due to the presence of high magnetic field strengths in $\sigma>>1$ regimes. We stress that a simple extrapolation and comparison of plasma parameters such as the magnetization from simulations to trace the edge or width of the jet observed at mm-wavelengths (cf. Lu et al. 2023, Fig. 3: grey band shows $\sigma$-contours from force-free electrodynamics simulations of a BZ jet for different BH spins) is insufficient to approximate the complex emission processes, that are yet to be fully understood. From the GRRT of our GRMHD simulations for the jet of M87* it is evident that the apparent width of the jet base is not anchored at the ergosphere of the BH, opposed to force-free electrodynamics simulations.
@@
Overall we find the jet of the K.99-model to exhibit more flux downstream the jet compared to the reference models. This aids in solving the issue that GRMHD simulations produce jets that decline too much in flux downstream compared to observations. Hence, this model is well suited for modelling jets on larger scales, beyond the jet launching region. This is very helpful, since most SMBH-systems in the universe are of smaller angular-size to us than M87*. With the enhanced jet power in this model, the jet remains brighter and more collimated on scales even beyond $1000\,\mathrm{M}$.
@@
The magnetic filling factor regulates the spatial extent of the intial vector potential and thus that of the magnetic field, as well as the total magnetic energy inside the initial torus configuration of our simulations, that are commonly used to model magnetically arrested accretion and also constitutes the main portion of the simulation library used for M87 by the EHT Collaboration et al.. In general, we conclude from this study that the evolution of these MAD simulations is very sensitive to minor changes in the initial magnetic field in the FM torus, set by the filling factor. In order to keep track of the vast parameter space of MAD simulations and consequences of the initial conditions, the specification of the detailed, initial vector-potential of simulations is therefore paramount.
@@
Acknowledgements. We thank Charles F. Gammie for the helpful discussion regarding this work. This research is supported by the DFG research grant “Jet physics on horizon scales and beyond” (Grant No. 443220636) within the DFG research unit “Relativistic Jets in Active Galaxies” (FOR 5195). The numerical simulations and calculations have been performed on MISTRAL at the Chair of Astronomy at the JMU Wuerzburg. YM is supported by the National Key R&D Program of China (grant no. 2023YFE0101200), the National Natural Science Foundation of China (grant no. 12273022), and the Shanghai municipality orientation program of basic research for international scientists (grant no. 22JC1410600)
@@
Appendix A Initial Magnetic Field Configuration and Filling-scaling of the Tori
@@
Here, we calculate the initial covariant volume fractions of the volume containing the cells in which the initial invariant magnetic field strength $b$ is defined and the initial volume enclosing the full mass density distribution of the torus. The initial magnetic field is normalized such that the minimum of plasma-$\beta$ in the torus is $\beta_{\rm{min}}=100$ in all simulations. We calculate the total and shell-integrated covariant volumes of the torus and enclosed magnetic field as follows:
$V_{\mathrm{t}}=\int_{\{rho>8\times 10^{-7}\}}\sqrt{-g}\,dr\,d\theta\,d\phi,\qquad$ $V_{\mathrm{b}}=\int_{\{b>1\times 10^{-7}\}}\sqrt{-g}\,dr\,d\theta\,d\phi,$ (31)
$V_{\mathrm{t}}(r)=\int_{\{rho>8\times 10^{-7}\}}\sqrt{-g}\,d\theta\,d\phi,\qquad$ $V_{\mathrm{b}}(r)=\int_{\{b>1\times 10^{-7}\}}\sqrt{-g}\,d\theta\,d\phi,$ (32)
We choose the masking conditions for the enclosed volume of the density distribution such that $\rho$ is above the density distribution in the atmosphere (the space excluding the torus and BH), that should have its minimum at $\rho_{\mathrm{min}}=10^{-6}$ (as stated in Sec. 2), however, increases up to $\rho\sim 8\times 10^{-7}$ in the vicinity of the BH. For $b$, the condition is not as strict, since the its numeric value is only defined, where the four-vector potential $A_{\phi}$ contributes to the curl to generate the magnetic field. We choose a low threshold of $b>1\times 10^{-7}$ and set the non-defined values that are mostly outside the torus to 0 in order to capture the full volume of the initial magnetic field. The actual filling factor is then given by:
$f_{\mathrm{a}}=\frac{V_{\mathrm{b}}}{V_{\mathrm{t}}},\quad f_{\mathrm{a}}(r)=\frac{V_{\mathrm{B}}(r)}{V_{\mathrm{torus}}(r)}.$ (33)
We note that, $f_{\rm{a}}$ is a meassure for the actual volumetric magnetic field extent (volumetric filling forth on), or the magnetic filling of the initial magnetospheric configuration, that we calculate for each filling factor $F$ and is not equivalent to the latter values as can be seen in Tab. 3.
@@
The total volumina contained by magnetic fields is generally quite close to the constant torus volumes throughout the simulations (cf. Fig. LABEL:fig:_initial_torus_filling), much closer than the filling factors $F$ suggest. However $F$, simply scales the offset of the initial vector potential $A_{\phi}$ cf. Eq. (6), essentially increasing its offset from the zero cut-off maximum function in Eq. (6) as can also be seen in Fig. 1. Still, increasing $F$ indeed leads to an increased volume that contains magnetic fields. Hence, we use the filling factor synonymously with the scaling of the magnetic field extent and the magnetic energy in the FM-torus.

### v2

@@ 3.3 Jet edge and width profiles
Figure 13: $\langle I\rangle=5\times 10^{-6}\,[\mathrm{Jy}]$ contours of time-averaged 86 GHz ray tracings (not the line integral convolved images shown in Fig. 11) for simulations K.80 (green), K.90 (blue), and K.99 (red). The colored areas are $1\sigma$ bands of the standard deviations in the toroidal direction.
Figure 14: Jet width of M87 at 86 GHz along the longitudinal jet axis for simulations K.80 (green), K.90 (blue), and K.99 (red) with corresponding $1\sigma$ bands. The black dots are data points of the jet width of M87 from the observation of Lu et al. (2023). The dotted line is a power-law fit from observations of M87. The solid horizontal line is the estimated 86 GHz ring size, and the dashed horizontal line is the 230 GHz ring size with respective error band (gray areas). The vertical lines are the distances from the BH, where the intrinsic half-opening angles equals the jet viewing angle of $\theta_{\mathrm{v}}=17^{\circ}$ (color-code is consistent).
From these jet profiles of the $86$ GHz ray-tracing we can now extract the jet widths as a function $W(z)$ of the jet longitudinal axis $z$ (as defined before) and compare these to the observations of Lu et al. (2023) shown in Fig. 14. We note that the data-points of the jet-width measurements do not show the error-bars in Lu et al. (2023), hence we cannot elaborate on the statistical certainty of this comparison. From Fig. 14, it is apparent that the theoretical predictions especially of K.80 and K.90 do match the observations well at the scales below $100\,\mu as$, contrarily to the claims of Lu et al. (2023) that MAD simulations may have difficulty explaining the observed jet width at these scales. At the scales shown above $100\,\mathrm{\mu as}$ the simulations K.80 and K.90 converge to the extrapolated $W(z)\propto z^{0.58}$ power law fit from large scale observations (see Asada and Nakamura (2012), Hada et al. (2013), Hada et al. (2016)). K.99 follows the power-law further up-stream and exhibits a sudden increase of the jet width at $\sim 40\,\mathrm{\mu as}$, whilst even closer to the BH it becomes almost constant, but even wider than the other models. This sudden increase in the width of the jet can be explained by the counter-jet, which exhibits a wider opening angle and thus shifts the jet edge of the counter-jet to positive $z$-values up to $\sim 100\,\mathrm{M}$ (cf. Fig. 11, right image). For the common choice of the R-$\beta$ electron temperature- and $\kappa$ non-thermal electron prescriptions chosen here K.99 exhibits a steeper slope than this observation suggests, rather following the large-scale power-law. The new observational data points at the scales closer to the jet launching region suggest that K.90 satisfies these observations very well, as the data points are very well within the variability of K.90 (cf. green band in Fig. 14). We also see that this holds mostly for the common model K.80, as well.
@@ 4 Conclusions
We find that the initial magnetic energy content of the torus systematically alters the jet energetics and morphology of the developed simulations. It mostly affects the EM contribution in the jet, and from that it gains $\sim 60\%$ efficiency in the conversion from rest-mass energy to the energy of the outflow by means of increasing the magnetic energy of the torus at the initialization. This in turn can make nonthermal particle acceleration from vacuum gaps and subsequent pair cascades more likely, due to the increased EM field in the jet spine (see Levinson and Rieger (2011), Wendel et al. (2021) for details on the vacuum-gap model and applications to low-luminosity AGNs such as M87 and Mrk501). We also find, that an increased magnetic filling of the FM torus elevates the outward angular-momentum transport, primarily in the jet, which is driven by Maxwell stresses. Comparing the angular-momentum transport of BHs with very high spin presented here to those of Chatterjee and Narayan (2022), in which the case of a non-rotating BH in a standard MAD-setup ($F=0.8$) was investigated, the major differences are the following: $i)$ the total radial angular-momentum flux is dominated by radially outward-directed flux in the jet and wind, mostly originating from the jet base; and ii)ii) morphologically, the angular-momentum vector fields are distinct, in that the high-spin BH creates field lines that point radially outwards in a polar angle range from $\sim(0-70)^{\circ}$ or $\sim(110-180)^{\circ}$. Toward the equatorial plane, the field lines transition from a monopole geometry to a quadrupole geometry in the disk and torus. The dominant, Maxwell-stress-induced outward polar flux is redirected to an equatorially inward, advectively driven flux. In contrast, the case of a non-rotating BH exhibits dominant, advectively driven inward angular-momentum flux in a larger portion of the disk and torus, whereas the outward-directed Maxwell-stress-induced flux is subdominant in the jet, such that the net angular momentum is transferred to the central BH. The outward angular momentum transport also covers a much smaller polar angle compared to simulations with high spin, contributing far less to the strength of outward winds. Our simulations of MADs with high-spin BHs demonstrate the effect of extracting the angular momentum of the BH to the angular momentum of the outflowing plasma in the jet and wind, as originally proposed by Blandford and Znajek (1977). We further find that there is a maximum of all net angular momentum fluxes inside the ergopsheric region of the BH, as opposed to the net-constant fluxes for MAD simulations with no spin.
@@
It appears that the disk’s vertical flux tubes created by the magnetically arrested disk affect the jet structure more significantly for high filling factors. From our sampled MAD states, we find a tendency for more spatially extended flux tubes that affect the jet morphology by introducing significant azimuthal asymmetries in the case of K.99; these propagate along the jet at least up to $200\,\mathrm{M}$. The jet’s sheath structure is more stable on these scales; i.e., the jet structure exhibits lower variability in the jet-perpendicular direction. However, since we restricted the jet-morphology analysis to a time interval, in that a single flux eruption occurs in each simulation, further detailed, statistical analyses of the impact of the filling factor on the quasi-periodic flux eruptions could provide valuable insights into the observable footprints that they have.
@@
The radiative transport calculations also show that the jet’s sheath structure is more stable and collimated up to $\sim 200\,\mu as,$ supported by the higher EM energy flux contribution in K.99. We find significantly elevated LP fractions of up to $\sim 15\%,$ in agreement with observations of the EHTC et al. (2021b). The jet width relation over the jet axial distance shows that GRRT calculations of GRMHD simulations of magnetically arrested accretion flow are able to match the observed jet width profile of Lu et al. (2023) very well for the reference models. For K.99, we find deviations from the jet width, as it follows the observed large-scale power-law relation $W\propto z^{0.58}$ to shorter distances, whereas the observations follow a sub-power-law relation. The deviation of K.99 from the observations, however, is a result of the different techniques used for the determination of the jet width. In the observations of Lu et al. (2023), the HWHM of the outermost components perpendicular to the jet longitudinal axis determines the jet width, whereas we did not recreate the edge brightened jet morphology and thus used brightness contours to determine the jet width. This made the detailed jet–width relation difficult to compare to these observations; nevertheless, we show that all our MAD simulations provide a wide jet base substantially wider than the pure Blandford–Znajek (BZ) jet (see Blandford and Znajek 1977) for a high-spin BH ($a\rightarrow 1$). It is evident from our analysis of the jet energy fluxes that the almost evacuated, highly magnetized jet spine ($\sigma\gtrsim 1$) is dominated by EM energy flux, and in the enveloping jet sheath the wind contributes to the outflowing kinetic energy flux that increases with distance. This jet-sheath region ($\sigma\lesssim 1$ and the gravitationally unbound part of the plasma) is the origin of the jet emission, and the $\sigma=1$ contour is thus not a good tracer for the outermost jet edge that could be observed. We remind the reader that $\sigma=\sigma_{\mathrm{cut}}=1$ is the value for the emission cutoff for the jet spine, where we do not expect significant synchrotron emission in the radio regime because of the extremely low electron densities, as well as efficient synchrotron cooling, due to the presence of high magnetic-field strengths in $\sigma>>1$ regimes. We stress that a simple extrapolation and comparison of plasma parameters such as the magnetization from simulations to trace the edge or width of the jet observed at millimeter wavelengths (see Lu et al. 2023, Fig. 3; gray band shows $\sigma$ contours from force-free electrodynamic simulations of a BZ jet for different BH spins) is insufficient to approximate the complex emission processes, which are yet to be fully understood. From the GRRT of our GRMHD simulations for the jet of M87*, it is evident that the apparent width of the jet base is not anchored at the ergosphere of the BH, as opposed to force-free electrodynamic simulations.
@@
Overall, we find the jet of the K.99-model to exhibit more flux downstream of the jet compared to the reference models. This aids in solving the issue that GRMHD simulations produce jets that decline too much in flux downstream compared to observations. Hence, this model is well suited for modeling jets on larger scales, beyond the jet launching region. This is very helpful, since most SMBH systems in the Universe are of smaller angular size than M87*. With the enhanced jet power in this model, the jet remains brighter and more collimated on scales even beyond $1000\,\mathrm{M}$.
@@
The magnetic filling factor regulates the spatial extent of the initial vector potential and thus that of the magnetic field, as well as the total magnetic energy inside the initial torus configuration of our simulations, which are commonly used to model magnetically arrested accretion and also constitute the main portion of the simulation library used for M87 by the EHTC et al.. In general, we conclude from this study that the evolution of these MAD simulations is very sensitive to minor changes in the initial magnetic field in the FM torus, which was set by the filling factor. In order to keep track of the vast parameter space of MAD simulations and consequences of the initial conditions, the specification of the detailed initial vector potential of simulations is therefore paramount.
@@
Acknowledgements. We thank Charles F. Gammie for the helpful discussion regarding this work. This research is supported by the DFG research grant “Jet physics on horizon scales and beyond” (Grant No. 443220636) within the DFG research unit “Relativistic Jets in Active Galaxies” (FOR 5195). The numerical simulations and calculations have been performed on MISTRAL at the Chair of Astronomy at the JMU Wuerzburg. YM is supported by the National Key R&D Program of China (grant no. 2023YFE0101200), the National Natural Science Foundation of China (grant no. 12273022, 12511540053), and the Shanghai municipality orientation program of basic research for international scientists (grant no. 22JC1410600)
@@
Appendix A Initial magnetic-field configuration and filling-scaling of the tori
@@
Here, we calculate the initial covariant volume fractions of the volume containing the cells in which the initial invariant magnetic-field strength $b$ is defined and the initial volume enclosing the full mass density distribution of the torus. The initial magnetic field is normalized such that the minimum of plasma $\beta$ in the torus is $\beta_{\rm{min}}=100$ in all simulations. We calculate the total and shell-integrated covariant volumes of the torus and enclosed magnetic-field as follows:
$V_{\mathrm{t}}=\int_{\{rho>8\times 10^{-7}\}}\sqrt{-g}\,dr\,d\theta\,d\phi,$ (31)
$V_{\mathrm{b}}=\int_{\{b>1\times 10^{-7}\}}\sqrt{-g}\,dr\,d\theta\,d\phi.$ (32)
We chose the masking conditions for the enclosed volume of the density distribution such that $\rho$ is above the density distribution in the atmosphere (the space excluding the torus and BH), that should have its minimum at $\rho_{\mathrm{min}}=10^{-6}$ (as stated in Sec. 2), however, increases up to $\rho\sim 8\times 10^{-7}$ in the vicinity of the BH. For $b$, the condition is not as strict, since its numeric value is only defined, where the four-vector potential $A_{\phi}$ contributes to the curl to generate the magnetic field. We choose a low threshold of $b>1\times 10^{-7}$ and set the non-defined values that are mostly outside the torus to 0 in order to capture the full volume of the initial magnetic field. The actual filling factor is then given by:
$f_{\mathrm{a}}=\frac{V_{\mathrm{b}}}{V_{\mathrm{t}}}.$ (33)
We note that, $f_{\rm{a}}$ is a measure for the volumetric magnetic-field extent, or the magnetic filling of the initial magnetospheric configuration, that we calculate for each filling factor $F$ and is not equivalent to the latter values as can be seen in Tab. 3.
@@
The total volume containing magnetic fields is generally quite close to the constant torus volumes throughout the simulations, much closer than the filling factors $F$ suggest. However $F$, simply scales the offset of the initial vector potential $A_{\phi}$ (cf. Eq. (6)), essentially increasing its offset from the zero-cutoff maximum function in Eq. (6) as can also be seen in Fig. 1. Still, increasing $F$ indeed leads to an increased volume that contains magnetic fields. Hence, we use the filling factor synonymously with the scaling of the magnetic-field extent and the magnetic energy in the FM-torus.